Suppose that X and Y are independent random variables with the same unknown mean u. Both X and Y have a variance of 36. Let T = aX + bY be an estimator of u. What condition must a and b satisfy in order that T be an unbiased estimator for ? Is T a normal random variable?
Answer #1

Given T=aX+bY

for T to be an unbiased estimator, Expectation(E) should be equal to the population mean u.

this means that

E(T)=aE(X)+bE(Y)

to be an unbiased estimator, E(X)=U and E(Y)=U

Thus U=aU+bU so the condition needed is that a+b should be equal to 1.

Yes, T is a normal random variable with a mean U and variance=a var(X)+b var(Y)=a(36)+b(36)=36 as a+b=1.